OpenStudy (anonymous):

Help please i will fan and medal What is the solution to the equation? (x + 8) - 14 = -15 x =_____

OpenStudy (johnweldon1993):

Our goal is to isolate 'x' So lets add 14 to both sides of this equation \[\large (x + 8) - 14 + 14 = -15 + 14\] \[\large (x + 8) = -1\] \[\large x + 8 = -1\] Now we just need to subtract 8 from both sides of the equation to solve for 'x' \[\large x = -1 - 8 \] \[\large x = ?\]

OpenStudy (anonymous):

Im probably wrong but -7 or 7

OpenStudy (johnweldon1993):

If it were 1..and not -1...then you would be correct :) but since we have -1... -1...minus another 8....would leave us at -9 right? :)

OpenStudy (anonymous):

Oh Yes! because there diffrent signs right?

OpenStudy (johnweldon1993):

Well, not entirely..just kinda think of it as a number line Where we start...is -1 and we need to move back 8 more